# Leveraging Prior Mean Models for Faster Bayesian Optimization of Particle Accelerators

### v3

@@ Introduction
Particle accelerators are complex machines that have thousands of free parameters which can be tuned to increase accelerator performance and beam quality. Controlling these parameters in real time during accelerator operations through the use of advanced algorithms, so-called online control, enables more complex beam manipulation capabilities and can help reduce the time spent on conducting routine tuning tasks. However, using optimization algorithms to solve facility-scale optimization problems that contain hundreds of free parameters is an open challenge due to the exponential scaling of parameter space volume with increasing numbers of parameters, often referred to as the “curse of dimensionality” [1]. When developing effective algorithms for accelerator control, it is important to incorporate prior information about beam physics into the algorithm to efficiently find solutions in extremely large parameter spaces. Additionally, many accelerator facilities frequently switch beam conditions for different users. This is the case, for example, at both the Linac Coherent Light Source (LCLS) and the Argonne Tandem Linear Accelerator System (ATLAS), a heavy ion accelerator. While these configurations can be treated as separate optimization problems, the time needed to switch between conditions could be reduced by leveraging information from related machine configurations.
@@
Bayesian optimization (BO) [2] is one such algorithm type that has been used in accelerators and can incorporate prior physics information into the decision-making process. The BO algorithm builds a Gaussian process model (GP) [3] that uses Bayesian statistics to predict objective function values (and corresponding uncertainties) from measurements of the objective function. This model is then used to inform control decisions in search of a globally optimal parameter set. Previous work [4, 5, 6] has demonstrated that including prior information from physics into BO algorithms increases their ability to efficiently converge to optimal solutions in high-dimensional parameter spaces. The use of Bayesian statistics (which inherently takes prior information into account) to build the GP model provides a natural framework for incorporating physics information into the modeling of accelerator-related optimization objectives. For example, GP models are built using a prior mean function, which specifies the prediction of the GP model in the absence of data. These prior mean functions are usually specified to be uninformative, as shown in Fig. 1(a). However, if prior information about the objective function is known, either from beam dynamics laws, simulations or prior data, an informative prior mean can be specified to incorporate this information directly into the GP model, as shown in Fig. 1(b). This can increase the accuracy of the GP model which helps guide the search for an optimum, improving the algorithm’s ability to converge to solutions quickly in high-dimensional spaces.
@@
In this work, we demonstrate and evaluate the use of fast-executing neural network (NN) surrogate models trained on beam dynamics simulations and historical data sets, as informative, physics-based prior mean functions in GP models to improve online accelerator control for two impactful use cases: injector tuning for the LCLS and tuning for different ion species at ATLAS. In particular, we demonstrate applications of this method in the context of realistic accelerator operations, including adapting surrogate models trained on beam dynamics simulations to online control at the LCLS injector, and transfer learning between different ion species at the ATLAS heavy ion accelerator, with improvements in convergence speed. These two cases were carefully chosen based on the expected benefit of the NN prior approach. The LCLS has to frequently switch setups for different users (different beam energies, bunch lengths, etc), and the injector is one of the more challenging parts of the machine to optimize due to space charge forces and the resulting sensitivity of the system to changes in controlled and uncontrolled inputs (such as drift in the initial laser spot). The injector is also not continuously tuned, because the relevant diagnostics are destructive. This makes it appealing both for using BO and for leveraging NN priors obtained from simulations and historical data. In addition, detailed physics simulations and ML-based surrogates of the LCLS injector are readily available. Similarly, ATLAS has to switch frequently between different ion species. Complicating this, beam dynamics for heavy ion linacs are very complicated and sensitive to machine inputs and changing conditions. Leveraging information across different configurations to speed up switching between ion species would substantially aid operation. The work presented here expands on previous work [7, 8, 9] that was conducted in a more limited scope. The first shows an offline proof-of-concept demonstration of using a NN prior for emittance and matching tuning on the LCLS injector. The second is a preliminary experimental demonstration at FRIB using an NN prior from historical data for tuning. The third is an initial experimental demonstration at ATLAS. We also explore the impact that the accuracy of the prior model has on the speed of convergence, and present a method to counteract the effects of an inaccurate model. As physics models often deviate substantially from real machine behavior, handling of machine-model mismatch is a major consideration for practical use of the NN prior approach. We show that combining approximate or well-calibrated NN models with GP models is an effective, computationally inexpensive way to incorporate detailed physics information into BO algorithms, significantly reducing the number of optimization steps needed to converge to a solution, even in cases where the NN model does not perfectly predict the experimental objective. Our demonstration also shows the utility of using prior information across ion species to improve tuning speed, which is important for current and future nuclear physics machines.
@@ Results
where and denote the beam size in their respective dimension. While the original model is used to represent the objective function , the prior mean function is given by variations of the same model with different levels of accuracy relative to the original. This is a more realistic scenario than the particular case of a perfect prior mean model, in which the optimization problem becomes trivial and is typically solved within a single step. The model accuracy is reduced by training a randomly initialized neural network with the same architecture as the LCLS surrogate on a grid of data points while limiting the number of training epochs. Models that are trained over fewer epochs show significant disagreement with respect to the LCLS surrogate, while models trained over more iterations tend to be more accurate.
@@
Based on the general shape of the performance curves in Fig. 3, we linearly decrease the weight from an initial to at step . As shown in Fig. 5 (green line), this can improve the robustness of Bayesian optimization with a non-constant prior mean, bringing performance closer to that with a constant prior. To take this idea one step further, we let the weighting factor be chosen based on the correlation coefficient determined across previous samples
@@
The LCLS surrogate model trained on data generated by the LCLS IMPACT-T model was used as an NN prior mean to perform online optimization of the LCLS photoinjector beamline. We repeatedly performed BO with the EI acquisition function starting with a fixed set of three randomly chosen initial points, with and without the NN prior mean function, to minimize the objective function in Eq. 1. During this, the laser spot size, pulse length, RF cavity parameters and bunch charge were fixed, resulting in a 9-dimensional optimization problem. Results from these optimization runs can be seen in Fig. 6. We observe that using the NN prior model consistently outperforms the constant prior in the first optimization steps. It takes the constant prior version five steps longer on average than the NN prior version to reach the same optimum. However, after approximately ten steps, the constant prior model surpasses the optimum achieved by the NN prior model, likely due to its low correlation with the measured objective (, ). The NN prior mean model provides an advantage in speed during the early stages of coarse optimization, and it is then surpassed in later refinement of the optimum by standard BO. Based on the simulation results discussed above, it is expected that NN models that are better calibrated to experimental data will increase the performance of BO in this context. In the meantime, it is possible to improve the optimization performance at later steps by transitioning from the NN model to a constant prior mean during the optimization, as demonstrated in the simulation section.
@@ Discussion
In this work we have experimentally demonstrated how incorporating a fast executing NN surrogate model of the target system can dramatically improve the performance of BO algorithms for two important use-cases: injector optimization at LCLS and switching between ion species at the ATLAS heavy ion linac. We also show in a simulation example that in cases where the prior model is more accurate and shows stronger correlation with ground truth values, the algorithm significantly outperforms standard BO with a constant prior mean function. These NN based prior mean models can be used to efficiently incorporate information from beam dynamics simulations, historical data, and similar accelerator operating configurations to accelerate online tuning of accelerators, all while quickly adapting to discrepancies between surrogate model predictions and observations on the real machine. However, care must be taken if the prior mean model is poorly correlated or anti-correlated with the true objective function, which can reduce BO performance below the performance with a constant prior. As was observed in the ATLAS experimental demonstration, the use of an expressive prior mean model can also facilitate transfer learning across different beam setups. This can aid rapid switching between beam setups at accelerators that serve a variety of experiments with different requirements (e.g. different species of beams as in the ATLAS case, or different electron beam characteristics requested by photon science users at the LCLS).
Incorporating NN based priors into GP modeling is also a path toward enabling facility-scale BO with hundreds to thousands of input parameters. Neural network surrogate models are able to ingest extremely large sets of historical and simulated data without suffering from an increase in prediction cost, unlike GP models, which scale poorly with the number of training points. By using NN based surrogate models as prior means we can incorporate much more information about the objective function into BO than we would be able to otherwise. Additionally, combining NN surrogates with GP modeling improves the applicability of using NNs in situations where the NN model does not perfectly predict the real objective function. Gaussian process models can quickly adapt predictions to new data, whereas NN surrogates need significant amounts of data and training time to be calibrated to experimental conditions. Combing the two model types leverages the advantages of both modeling strategies and can be used to provide accurate, facility-scale models of objective functions that are updated in real-time.
@@ A Gaussian process [3] (GP) is a distribution of functions denoted as
that is fully characterized by its mean function and covariance or kernel function . It is a generalization of the multivariate Gaussian distribution for which each finite collection of function values follows a joint multivariate normal distribution.
In applications of BO and to simplify calculations, the prior mean function is commonly specified as . Given a set of collected data samples , the posterior distribution evaluated on test points can then be expressed as
@@
where denotes the covariance matrix between the respective sets of data and is the identity matrix. Note that to account for a constant mean function that is not zero, i.e., , we can simply redefine the objective function as
@@
The GP model is then passed to an acquisition function, which uses the model to predict the value of potential future measurements towards finding a global extremum. Popular acquisition functions, such as Expected Improvement (EI) and Upper Confidence Bound (UCB) [17], strike a balance between choosing points that add more information to the global GP model (exploration) and choosing points that are near predicted extremum (exploitation). Points in parameter space that maximize the acquisition function are then observed experimentally and added to the model data set. This process repeats sequentially until a convergence criteria is met. Regardless of the choice of acquisition function, the performance of BO algorithms depends on accurate GP surrogate modeling of the objective function using as few measurements as possible, which we examine in this work.
@@
We use two metrics to quantify the quality of the NN prior model in the context of improving BO performance. While a simple mean squared error (MSE) loss is used in training the NN parameters, this is an insufficient metric to assess the overall model quality for its use as a prior mean function in BO, as illustrated in Fig. 11. We see in Fig. 11(a) that using a prior mean model which has a relatively low MSE, but is poorly correlated with the ground truth objective function, will predict optimal locations that do not correspond to real extrema in the ground truth. However, if the prior mean model strongly correlates with the ground truth function, as shown in Fig. 11(b), predicted optimal locations agree well with the real optimal locations, even if the MSE of the model is larger. In our studies, we use the correlation coefficient
where is the Pearson correlation coefficient for a collection of samples, which is typically a more accurate metric in assessing model quality and predicting BO performance. Furthermore, as a slightly more intuitive measure of the absolute error than MSE, we augment the correlation coefficient with the mean absolute error (MAE) to describe the overall model quality.
@@ Acknowledgments
Conceptualization, T.B., J.L.M., B.M., D.R., R.R., and A.L.E.; Data curation, T.B., J.L.M., C.X., and K.R.L.B.; Formal analysis, T.B. and J.L.M.; Funding acquisition, D.R., A.L.E., and B.M.; Investigation, T.B. and J.L.M.; Methodology, R.R., T.B., J.L.M., and A.L.E.; Software, R.R., C.X., K.R.L.B., T.B., and J.L.M.; Experiment, T.B., J.L.M., R.R., and A.L.E.; Supervision, D.R., A.L.E., R.R., and B.M.; Validation, T.B. and J.L.M.; Visualization, T.B. and J.L.M.; Writing—original draft, T.B., J.L.M., R.R., B.M., and A.L.E.; Writing—review and editing, T.B., J.L.M., R.R., D.R., B.M., and A.L.E. All authors have read and agreed to the published version of the manuscript.

### v4

@@ Introduction
Particle accelerators are complex machines that have hundreds of free parameters which can be tuned to increase accelerator performance and beam quality. Controlling these parameters in real time during accelerator operations through the use of advanced algorithms, so-called online control, enables more complex beam manipulation capabilities and can help reduce the time spent on conducting routine tuning tasks. However, using optimization algorithms to solve facility-scale optimization problems that contain hundreds of free parameters is an open challenge due to the exponential scaling of parameter space volume with increasing numbers of parameters, often referred to as the “curse of dimensionality” [1]. When developing effective algorithms for accelerator control, it is important to incorporate prior information about beam physics into the algorithm to efficiently find solutions in extremely large parameter spaces. Additionally, many accelerator facilities frequently switch beam conditions for different users. This is the case, for example, at both the Linac Coherent Light Source (LCLS) and the Argonne Tandem Linear Accelerator System (ATLAS), a heavy ion accelerator. Injector tuning at LCLS by hand can often take hours, both for restoring the beam after a shut-down or when switching between different beam conditions (e.g. adjusting the injector for different beam charge setups). While these configurations can be treated as separate optimization problems, the time needed to switch between conditions could be reduced by leveraging information from related machine configurations.
Bayesian optimization (BO) is one such algorithm type that has been used in accelerators and can incorporate prior physics information into the decision-making process [2]. BO methods have been shown to reduce the tuning time in several applications when compared to more traditional methods (e.g. see [2, 3, 4]). The BO algorithm builds a Gaussian process model (GP) [5] that uses Bayesian statistics to predict objective function values (and corresponding uncertainties) from measurements, while factoring in a prior belief about the objective function. This model is then passed to an acquisition function to decide the next point to be observed in search of a globally optimal parameter set. Popular acquisition functions, such as Expected Improvement (EI) and Upper Confidence Bound (UCB) [6], strike a balance between choosing points that add more information to the global GP model (exploration) and choosing points that are near predicted extremum (exploitation).
@@
Previous work [4, 3, 7] has demonstrated that including prior information from physics into BO algorithms increases their ability to efficiently converge to optimal solutions in high-dimensional parameter spaces. The use of Bayesian statistics (which inherently takes prior information into account) to build the GP model provides a natural framework for incorporating physics information into the modeling of accelerator-related optimization objectives. For example, GP models are built using a prior mean function, which specifies the prediction of the GP model in the absence of data. These prior mean functions are usually specified to be uninformative, as shown in Fig. 1(a). However, if prior information about the objective function is known, either from beam dynamics laws, simulations or prior data, an informative prior mean can be specified to incorporate this information directly into the GP model, as shown in Fig. 1(b). This can increase the accuracy of the GP model which helps guide the search for an optimum, improving the algorithm’s ability to converge to solutions quickly in high-dimensional spaces.
@@
In this work, we demonstrate and evaluate the use of fast-executing neural network (NN) surrogate models trained on beam dynamics simulations and historical data sets, as informative, physics-based prior mean functions in GP models to improve online accelerator control for two impactful use cases: injector tuning for the LCLS and tuning for different ion species at ATLAS. In particular, we demonstrate applications of this method in the context of realistic accelerator operations, including adapting surrogate models trained on beam dynamics simulations to online control at the LCLS injector, and transfer learning between different ion species at the ATLAS heavy ion accelerator, with improvements in convergence speed. These two cases were carefully chosen based on the expected benefit of the NN prior approach. The LCLS has to frequently switch setups for different users (different beam energies, bunch lengths, etc), and the injector is one of the more challenging parts of the machine to optimize due to space charge forces and the resulting sensitivity of the system to changes in controlled and uncontrolled inputs (such as drift in the initial laser spot). The injector is also not continuously tuned, because the relevant diagnostics are destructive. This makes it appealing both for using BO and for leveraging NN priors obtained from simulations and historical data. In addition, detailed physics simulations and ML-based surrogates of the LCLS injector are readily available. Similarly, ATLAS has to switch frequently between different ion species. Complicating this, beam dynamics for heavy ion linacs are very complicated and sensitive to machine inputs and changing conditions. Leveraging information across different configurations to speed up switching between ion species would substantially aid operation. The work presented here expands on previous work [8, 9, 10] that was conducted in a more limited scope. The first shows an offline proof-of-concept demonstration of using a NN prior for emittance and matching tuning on the LCLS injector. The second is a preliminary experimental demonstration at FRIB using an NN prior from historical data for tuning. The third is an initial experimental demonstration at ATLAS. We also explore the impact that the accuracy of the prior model has on the speed of convergence, and present a method to counteract the effects of an inaccurate model. As physics models often deviate substantially from real machine behavior, handling of machine-model mismatch is a major consideration for practical use of the NN prior approach. We show that combining approximate or well-calibrated NN models with GP models is an effective, computationally inexpensive way to incorporate detailed physics information into BO algorithms, significantly reducing the number of optimization steps needed to converge to a solution, even in cases where the NN model does not perfectly predict the experimental objective. We also examine the impact of different levels of mismatch between the NN model and the accelerator on optimization speed, which is an important consideration. We find a high degree of model accuracy is not required; even coarse models still provide a benefit to BO convergence speed. Our demonstration also shows the utility of using prior information across ion species to improve tuning speed, which is important for current and future nuclear physics machines.
@@ Results
where and denote the beam size in their respective dimension. While the original model is used to represent the objective function , the prior mean function is given by variations of the same model with different levels of accuracy relative to the original. This is a more realistic scenario than the particular case of a perfect prior mean model, i.e., a model that perfectly predicts the objective function . In the latter case, the optimization problem becomes trivial and is typically solved within a single step. The model accuracy is reduced by training a randomly initialized neural network with the same architecture as the LCLS surrogate on a grid of data points while limiting the number of training epochs. Models that are trained over fewer epochs show significant disagreement with respect to the LCLS surrogate, while models trained over more iterations tend to be more accurate. We generally use the Pearson correlation coefficient and the mean absolute error (MAE), which are detailed in the Methods section, to describe the model accuracy.
@@
Scaling up to High-Dimensional Problems
The performance benefits of using a NN prior mean model, as observed in Fig. 3, stem from the fact that an informative prior helps to guide the algorithm to the most relevant areas of the parameter space without as much initial exploration. This additional guidance of the BO sampling sequence is especially interesting, and most impactful, when the optimization is performed over very large parameter spaces where an uninformed naive search can take prohibitively long to reach satisfactory results. This effect is demonstrated in a larger simulation of the LCLS beamline where the machine is parametrized by a total of quadrupoles and the objective is the Free-Electron Laser (FEL) pulse energy. The NN prior mean model is trained on data from the beam optics tracking simulations based on Cheetah [14] and a FEL surrogate model developed using the simulation results from GENESIS [15]. We then run BO with different models as prior mean functions to optimize over input parameters while evaluating against the beam physics model based on Cheetah and the FEL surrogate. The results shown in Fig. 5 again demonstrate the improved performance that can be obtained by using an informative prior mean model, especially a high-correlation model that can facilitate more effective navigation of the high-dimensional parameter space. In contrast, standard BO takes significantly longer to converge to less satisfactory values while the NN prior based approach gets close to the optimum after only a few steps. To see how long it would take regular BO to converge, we ran it on the same problem. Even after steps, it has not yet reached the same performance the prior mean BO had after one initial step.
@@
Based on the general shape of the performance curves in Fig. 3, we linearly decrease the weight from an initial to at step . As shown in Fig. 6 (green line), this can improve the robustness of Bayesian optimization with a non-constant prior mean, bringing performance closer to that with a constant prior. To take this idea one step further, we let the weighting factor be chosen based on the Pearson correlation coefficient determined across previous samples (see Methods section)
@@
The LCLS surrogate model trained on data generated by the LCLS IMPACT-T model was used as an NN prior mean to perform online optimization of the LCLS photoinjector beamline. We repeatedly performed BO with the EI acquisition function starting with a fixed set of three randomly chosen initial points, with and without the NN prior mean function, to minimize the objective function in Eq. 1. During this, the laser spot size, pulse length, RF cavity parameters and bunch charge were fixed, resulting in a 9-dimensional optimization problem. Results from these optimization runs can be seen in Fig. 7. We observe that using the NN prior model consistently outperforms the constant prior in the first optimization steps. It takes the constant prior version five steps longer on average than the NN prior version to reach the same optimum. However, after approximately ten steps, the constant prior model surpasses the optimum achieved by the NN prior model, likely due to its low correlation with the measured objective, , , where denotes the Pearson correlation coefficient and MAE the mean absolute error as detailed in the Methods section. The NN prior mean model provides an advantage in speed during the early stages of coarse optimization, and it is then surpassed in later refinement of the optimum by standard BO. Based on the simulation results discussed above, it is expected that NN models that are better calibrated to experimental data will increase the performance of BO in this context. In the meantime, it is possible to improve the optimization performance at later steps by transitioning from the NN model to a constant prior mean during the optimization, as demonstrated in the simulation section.
@@ Discussion
In this work we have experimentally demonstrated how incorporating a fast executing NN surrogate model of the target system can dramatically improve the performance of BO algorithms for two important use-cases: injector optimization at LCLS and switching between ion species at the ATLAS heavy ion linac. We also show in a simulation example that in cases where the prior model is more accurate and shows stronger correlation with ground truth values, the algorithm significantly outperforms standard BO with a constant prior mean function. These NN based prior mean models can be used to efficiently incorporate information from beam dynamics simulations, historical data, and similar accelerator operating configurations to accelerate online tuning of accelerators, all while quickly adapting to discrepancies between surrogate model predictions and observations on the real machine. However, care must be taken if the prior mean model is poorly correlated or anti-correlated with the true objective function, which can reduce BO performance below the performance with a constant prior. As was observed in the ATLAS experimental demonstration, the use of an expressive prior mean model can also facilitate transfer learning across different beam setups. This can aid rapid switching between beam setups at accelerators that serve a variety of experiments with different requirements (e.g., different species of beams as in the ATLAS case, or different electron beam characteristics requested by photon science users at the LCLS).
Incorporating NN based priors into GP modeling is also a path toward enabling facility-scale BO with hundreds of input parameters. Although these large parameter spaces remain challenging to navigate, including a prior mean model can reduce the complexity by guiding the search to the most relevant areas. In very high-dimensional parameter spaces, the implicit optimization of the acquisition function may eventually also become a non-trivial task.
Neural network surrogate models are able to ingest extremely large sets of historical and simulated data without suffering from an increase in prediction cost, unlike GP models, which scale poorly with the number of training points. By using NN based surrogate models as prior means we can incorporate much more information about the objective function into BO than we would be able to otherwise. Additionally, combining NN surrogates with GP modeling improves the applicability of using NNs in situations where the NN model does not perfectly predict the real objective function. Gaussian process models can quickly adapt predictions to new data, whereas NN surrogates need significant amounts of data and training time to be calibrated to experimental conditions. Combining the two model types leverages the advantages of both modeling strategies and can be used to provide accurate, facility-scale models of objective functions that are updated in real-time.
@@ A Gaussian process [5] (GP) is a distribution of functions denoted as
that is fully characterized by its mean function and covariance or kernel function . It is a generalization of the multivariate Gaussian distribution for which each finite collection of function values follows a joint multivariate normal distribution. In applications of BO and to simplify calculations, the prior mean function is commonly specified as a zero mean, i.e., a function that is zero everywhere in the parameter space
Given a set of collected data samples , the posterior distribution evaluated on test points can then be expressed as
@@
where denotes the covariance matrix between the respective sets of data and is the identity matrix. Note that to account for a constant mean function that is not zero, i.e., a function which has a constant, non-zero value everywhere in the parameter space
we can simply redefine the objective function as
@@
The GP model is then passed to an acquisition function, which uses the model to predict the value of potential future measurements towards finding a global extremum. Points in parameter space that maximize the acquisition function are then observed experimentally and added to the model data set. This process repeats sequentially until a convergence criteria is met. Regardless of the choice of acquisition function, the performance of BO algorithms depends on accurate GP surrogate modeling of the objective function using as few measurements as possible, which we examine in this work.
@@
We use two metrics to quantify the quality of the NN prior model in the context of improving BO performance. While a simple mean squared error (MSE) loss is used in training the NN parameters, this is an insufficient metric to assess the overall model quality for its use as a prior mean function in BO, as illustrated in Fig. 12. We see in Fig. 12(a) that using a prior mean model which has a relatively low MSE, but is poorly correlated with the ground truth objective function, will predict optimal locations that do not correspond to real extrema in the ground truth. However, if the prior mean model strongly correlates with the ground truth function, as shown in Fig. 12(b), predicted optimal locations agree well with the real optimal locations, even if the MSE of the model is larger. In our studies, we use the Pearson correlation coefficient
where is the covariance and and denote the respective standard deviations. We find this is typically a more accurate metric in assessing model quality and predicting BO performance. This can for example be seen in Fig. 5, where the model with the highest correlation outperforms other versions, although their absolute error is smaller. Furthermore, as a slightly more intuitive measure of the absolute error than MSE, we augment the correlation coefficient with the mean absolute error (MAE)
to describe the overall model quality.
@@ Acknowledgments
Conceptualization, T.B., J.L.M., B.M., D.R., R.R., and A.L.E.; Data curation, T.B., J.L.M., C.X., K.R.L.B., and Z.Z.; Formal analysis, T.B. and J.L.M.; Funding acquisition, D.R., A.L.E., and B.M.; Investigation, T.B. and J.L.M.; Methodology, R.R., T.B., J.L.M., and A.L.E.; Software, R.R., C.X., K.R.L.B., J.M., T.B., and J.L.M.; Experiment, T.B., J.L.M., R.R., and A.L.E.; Supervision, D.R., A.L.E., R.R., and B.M.; Validation, T.B. and J.L.M.; Visualization, T.B. and J.L.M.; Writing—original draft, T.B., J.L.M., R.R., B.M., and A.L.E.; Writing—review and editing, T.B., J.L.M., R.R., D.R., B.M., and A.L.E. All authors have read and agreed to the published version of the manuscript.